The product of two consecutive integers is 8 less than 8 times their sum . Find the integers.
Answer by unlockmath(1688)

Hello,
Let's represent x as one integer and x+1 as the next integer. Now we can set up and equation.
x(x+1)=[x+(x+1)]8-8 Rewritten it will look like:
x^2+x=16x+8-8 Let's put all this on one side of the "=" so subtract 16x from both sides to result in:
x^2-15x=0 We can solve x by using the quadratic formula.
{{x = (-b +- sqrt( b^2-4*a*c ))/(2*a) }}
Plugging in the numbers gives us:
x = 15
So the 2 consecutive numbers are:
15 & 16
Check it out in the above equation.
RJ
Hey, Check out a new book I wrote at:
www.math-unlock.com